Use Substitution to solve each system of equations. If the system does not have exactly one solution, state if it has no solution or infinitely many solutions. .3x-.2y=.5 and x-2y=-5 thank you
Answer by stanbon(75887)

.3x-.2y=.5
x-2y =-5
----------------
Modify:
3x - 2y = 5
x - 2y = -5
----------------
Subtract 2nd from 1st to get:
2x = 10
x = 5
-------------------
Substitute in x - 2y = -5 to solve for "y":
5 - 2y = -5
-2y = -10
y = 5
---------------------
Ans: x = 5 ; y = 5
===========================
Cheers,
Stan H.